A cement truck delivers mixed cement to a large construction site. Let x represent the cycle time in minutes for the truck to leave the construction site, go back to the cement plant, fill up, and return to the construction site with another load of cement. From past experiences, it is known that the mean cycle time is μ = 46 minutes with σ = 9 minutes. The x distribution is approximately normal. What is the probability that the cycle time will exceed 61 minutes, given that it has exceeded 44 minutes?

a. 0.081

b. 0.588

c. 0.048

d. 0.588

e. 0.048
Answer #1

Ans:

P(X>61/X>44)=P(X>61 and X>44)

=P(X>61)/P(X>44)

     

z(61)=(61-46)/9=1.667

z(44)=(44-46)/9=-0.222

P(z>1.667)=0.0478

P(z>-0.222)=0.5879

P(z>1.667)/P(z>-0.222)=0.0478/0.5879=0.081

Option a is correct(0.081)